Conduct a cost analysis of gas versus electric vehicle. Find the break-even point assuming gasoline is $3.00/gallon with an average of 22 miles per gallon that costs $25,000. Compared to an electric vehicle for $35,000 that holds 250 miles per charge for $8.00. After how many miles will you need to drive to make the total cost to operate of the two cars the same? For simplicity, you may assume that regular maintenance is the same for the two vehicles.
Answer #1

Break-even point will be the number of miles traveled at which total cost under both options ( i.e. electric vehicle vs gasoline vehicle ) will be the same.

Let the required number of miles traveled = N

Total cost = Cost of the vehicle + Cost of fuel/electric charge

For electric vehicle:

Cost of vehicle = $35,000

Cost of electric charge for N miles = $8/250 x N ( since one charge at a cost of $8 runs the vehicle for 250 miles )

Total cost ($) = 35000 + 8.N/250

For gasoline vehicle :

Cost of vehicle = $25,000

Cost of gasoline = $3/22 x N ( since the vehicle records average 22 miles per Gallon)

Total cost($) = $25000 + 3.N/22

At break-even point ( i.e. after running for N miles), total cost under both the above options should be the same.

Therefore,

35,000 + 8N/250 = 25000 + 3N/22

Or, 3N/22 – 8N/250 = 35000 – 25000

Or,( 375N – 88.N) / 2750 = 10,000

Or, 287.N/2750 = 10,000

Or, N = 10,000 X 2750 /287

Or,N = 95818.81

Therefore, after 95818.81 miles of driving, the total cost to operate both the cars will be the same.